OpenStudy (anonymous):

Two cards are drawn from a fifty-two-card deck. What is the probability that the draw will yield an ace and a face card?

OpenStudy (anonymous):

I havent worked with probability in like a year XD i forgot how to :P

OpenStudy (anonymous):

Think of it like this. There are 4 face/ace cards to every suit.

OpenStudy (anonymous):

That is true

OpenStudy (anonymous):

4 cards times 4 gives us 16. There are 16 out of 56 cards that are face cards.

OpenStudy (anonymous):

I mean 52.

OpenStudy (anonymous):

yeah I know

OpenStudy (anonymous):

there are a couple ways to do this but first we need to know does it mean "and ace and then a face card" which is easy or does it mean "one of the two cards is and ace, the other is a face card" which is somewhat harder

OpenStudy (anonymous):

Sorry ^^''.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

And true, @satellite73.

OpenStudy (anonymous):

But if he meant the way I did it, then its a simple 16:52

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

But its not done yet. Find the common factor.

OpenStudy (anonymous):

Did I mess up somewhere? o.o''

OpenStudy (anonymous):

first card is an ace \[\frac{4}{52}\] second card is a face card given that the first card is an ace \[\frac{12}{51}\] AND \[\frac{4}{52}\times \frac{12}{51}\]

OpenStudy (anonymous):

that is "first ace, then face card" which might not be what the question is asking

OpenStudy (anonymous):

one is a face card, the other is an ace \[\frac{4\times 12}{\binom{52}{2}}\]

OpenStudy (anonymous):

There is only an ace and a face card. I just asked my teacher

OpenStudy (anonymous):

Only one ace and one face?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(\binom{52}{2}=\frac{52\times 51}{2}=1326\) your answer is \[\frac{4\times 12}{1326}\]

OpenStudy (anonymous):

Alright Satellite how did you get that answer

OpenStudy (anonymous):

the denominator is all the ways you can choose two cards out of fifty two i.e. "fifty two choose 2" written as \(\binom{52}{2}\) or sometimes \(_{52}C_2\) or even \(^{52}C_2\) and computed via \[\frac{52\times 51}{2}\]

OpenStudy (anonymous):

the numerator is the number of ways you can pick one ace out of four, which is \(4\) times the number of ways you can pick 1 facecard out of 12, which is 12

OpenStudy (anonymous):

Like @satellite73 said, \[P(1\text{ ace}\cap1\text{ face})=\frac{P(1\text{ ace})\cdot P(1\text{ face})}{P(2\text{ cards from deck}}=\frac{\binom 41\binom{12}1}{\binom{52}2}\]

OpenStudy (anonymous):

thanks